OpenStudy (anonymous):

All of the following equations have the same solution except _____. 3m + 2 = -10 2 - 3m = -10 -10 = -3m + 2 -3m + 2 = -10 Please explane so i counld know how you got the answer.

OpenStudy (compassionate):

Hi, do you know how to do this the quick and easy way? Do you have a wild guess?

OpenStudy (anonymous):

I am guessing it is the seconds one

OpenStudy (compassionate):

2 - 3m = -10 -3m = -12 m = 4 2 - 3m = -10 2 - 3(4) = -10 2 - 16 = 10 14 = 10 wrong

OpenStudy (anonymous):

then my second guess would be the first one

OpenStudy (compassionate):

Stop guessing and show some work

OpenStudy (anonymous):

It is the first one because it is -4 and the other ones are 4

OpenStudy (anonymous):

I did -10-2=-12 -12/3 which equals -4 Am i correct?